# Enhancing Generative Auto-bidding with Offline Reward Evaluation and Policy Search

### v2

@@ Abstract
Auto-bidding serves as a critical tool for advertisers to improve their advertising performance. Recent progress has demonstrated that AI-Generated Bidding (AIGB), which learns a conditional generative planner from offline data, achieves superior performance compared to typical offline reinforcement learning (RL)-based auto-bidding methods. However, existing AIGB methods still face a performance bottleneck due to their inherent inability to explore beyond the static offline dataset. To address this, we propose AIGB-Pearl (Planning with EvaluAtor via RL), a novel method that integrates generative planning and policy optimization. The core of AIGB-Pearl lies in constructing a trajectory evaluator for scoring generation quality and designing a provably sound KL-Lipschitz-constrained score maximization scheme to ensure safe and efficient generalization beyond the offline dataset. A practical algorithm incorporating the synchronous coupling technique is further devised to ensure the model regularity required by the proposed scheme. Extensive experiments on both simulated and real-world advertising systems demonstrate the state-of-the-art performance of our approach.
@@ 1 Introduction
Hence, there arises a question: built on AIGB, the latest state-of-the-art auto-bidding method, can we devise a plausible scheme to involve policy optimization in its generative model? To this end, a natural idea is to integrate offline RL methods into AIGB. However, it is nontrivial to implement in the auto-bidding problem since (i) there is a lack of reward signals in AIGB to guide the generative model. Specifically, the generation quality of the generative model remains unknown during training, making it infeasible to explore new trajectories beyond the offline dataset; (ii) no dedicated offline RL algorithm exists for AIGB. In particular, theoretical analysis that guarantees safe generalization and mitigates OOD issues for generative models in auto-bidding remains largely unexplored.
To address these critical challenges, we propose AIGB-Pearl (Planning with EvaluAtor via RL), an RL-enhanced version of AIGB that learns a trajectory evaluator to score generation quality and drive exploration through continuous interaction with the generative model in AIGB via RL. The evaluator is trained through supervised learning, leveraging the offline dataset. Notably, to mitigate the OOD problem and ensure reliable usage of the evaluator, we first examine the theoretical upper bound on the evaluator’s bias. Then, guided by this analysis, we establish a KL-Lipschitz-constrained score maximization objective with a provable sub-optimality bound, enabling safe and effective exploration for the generative model beyond the offline data. Moreover, to perform constrained score maximization, we design a practical algorithm incorporating the synchronous coupling technique, which helps ensure the Lipschitz requirement of the generative model. Besides, we note that AIGB-Pearl operates without the need for bootstrapping, leading to greater training stability compared to offline RL methods.
To summarize, our contributions in this paper are fourfold: (i) we propose a novel generative auto-bidding method, AIGB-Pearl, that enables continuous improvement in generation quality through exploration beyond the offline dataset; (ii) we propose a provable KL-Lipschitz constrained score maximization objective with a sub-optimality bound, ensuring a safe and effective generalization for AIGB; (iii) we devise a practical algorithm with synchronous coupling that effectively ensures the Lipschitz requirement for the generative model; (iv) extensive simulated and real-world experiments demonstrate that AIGB-Pearl achieves SOTA performance and verify the effectiveness of the developed techniques in enhancing safe and effective generalization.
@@ 2.2 Offline RL Methods
Due to the offline setting of the considered auto-bidding problem, directly employing Eq. 2 results in the infamous out-of-distribution (OOD) problem (Fujimoto et al., 2019), making the policy erroneously deviate from the offline dataset $\mathcal{D}$. As a standard solution, offline RL (Yu et al., 2020; Kumar et al., 2020; Kidambi et al., 2020; Wang et al., 2022) constrains the policy’s behavior near $\mathcal{D}$ during TD learning, enabling reliable generalization beyond the offline dataset.
@@ 2.3 Generative Auto-bidding Methods
Unlike RL methods, the AI-generated auto-bidding (AIGB) (Guo et al., 2024) treats the auto-bidding problem as a sequence generation task. Specifically, a generative model is employed to fit the conditional trajectory distribution $p_{\theta}(\tau|y(\tau))$ within the offline dataset $\mathcal{D}$, i.e.,
@@
where $\theta$ denotes the parameter. Let $y_{m}>0$ be the maximum trajectory quality in $\mathcal{D}$, we have $\forall y\in\mathcal{D},y\in[0,y_{m}]$. During inference, AIGB follows a planning-and-control architecture. Specifically, at each time step, a trajectory is sampled from the trained generative model that acts as the planner, with a manually set condition $y^{*}\triangleq(1+\epsilon)y_{m}$, where $\epsilon>0$ is a hyper-parameter with usually a small value. Then, an extra off-the-shelf inverse dynamic model (Agrawal et al., 2016), acting as the controller, is employed to compute the action. See Appendix B for detailed descriptions. AIGB avoids TD learning and generally outperforms offline RL methods (Guo et al., 2024).
@@ 3 Method
Enabling AIGB to explore higher-quality trajectories actively beyond the offline dataset can enhance its performance and generalization ability. To this end, we propose AIGB-Pearl (Planning with EvaluAtor via RL) that constructs a trajectory evaluator (referred to as the evaluator for simplicity) to integrate RL methods into AIGB’s planner training. Specifically, the evaluator learns a score $\hat{y}_{\phi}(\tau)$ to estimate the trajectory quality $y(\tau)$ via supervised learning, where $\phi$ denotes the evaluator parameter. Then, with $\phi$ fixed, the planner tries to maximize the score of its generation through constant interactions with the evaluator, as shown in Fig. 1. Formally, this can be formulated as:
@@
The effectiveness of AIGB-Pearl fundamentally hinges on the evaluator’s reliability. However, given the offline setting of the considered auto-bidding problem, the evaluator learning is confined to the offline dataset $\mathcal{D}$, and directly solving Eq. 4 can induce the infamous OOD problem due to the evaluator’s generalization limits, potentially resulting in the planner’s true performance degradation.
To address this challenge, we begin by examining the theoretical bounds on the evaluator’s bias. Then, guided by this analysis, we propose a KL-Lipschitz-constrained score maximization objective for the planner to ensure reliable evaluator utilization in Section 3.1. This objective is theoretically justified by a sub-optimality bound, as established in Section 3.1.1. Finally, a practical algorithm is presented in Section 3.2, where a synchronous coupling method is employed to realize the planner’s Lipschitz constraint.
@@ 3.1 KL-Lipschitz-constrained score maximization
Our basic idea is to optimize $\theta$ within a domain where the gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)\triangleq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]$ is bounded by a small certifiable upper bound. This ensures the score maximization occurs only in regions where the evaluator is reliable.
@@
Based on this property, we incorporate a $\sqrt{T}R_{m}$-Lipschitz continuity constraint into the evaluator’s design (as described in Section 3.2.1), and denote the Lipschitz constant of the resulting evaluator as $k\sqrt{T}R_{m}$, where $k\geq 0$ reflects the violation degree. A tighter satisfaction of the Lipschitz constraint by the evaluator results in $k$ closer to $1$.
@@
Note that $\delta_{D}$ could be regulated to a small value via supervised training of the evaluator. Consequently, bounding the planner’s performance gap requires constraining the following two factors:
@@
The proof is given in Appendix C.3. Therefore, we constrain the planner’s Lipschitz constant to a positive value $L_{p}>0$ to ensure the boundedness of the first Wasserstein term, where $L_{p}$ is a hyperparameter whose recommended range is provided later in Eq. 10.
@@
Intuitively, the KL and Lipschitz constraints collectively ensure the planner’s generation under condition $y^{*}$ remains within a certified neighborhood of the offline dataset $\mathcal{D}$, while the evaluator maintains high accuracy within this $\mathcal{D}$-proximal region as it is trained with the offline dataset. Therefore, Eq. 8 can realize reliable evaluator-guided score maximization.
@@
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the expected distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$. The true performance gap between the optimal parameter $\theta^{*}$ and the solution $\hat{\theta}$ to Eq. 8 is bounded by:
@@
Theoretical Result Analysis. In Theorem 3, the constants $\delta_{M},R_{m},T,y_{m},\epsilon,\rho$ characterize domain-specific properties of the auto-bidding task and the offline dataset $\mathcal{D}$. Nonetheless, a lower training error $\delta_{D}$ of the evaluator and a closer $k$ to $1$ correspond to a smaller sub-optimality gap. Note that $k$ cannot approach zero without compromising $\delta_{D}$, as excessively small $k$ prevents the evaluator from fitting the offline dataset $\mathcal{D}$.
@@ 3.2.1 Lipschitz Trajectory Evaluator
where $\beta_{1}>0$ is a hyper-parameter, $\cdot]_{+}\triangleq\max\{0,\cdot\}$.
@@ 3.2.2 Lipschitz Planner With Synchronous Coupling
Wasserstein Upper Bound as Surrogate. Accurate computation of this Wasserstein distance term is challenging, as it requires finding the optimal coupling between $p_{\theta}(\tau|y_{1})$ and $p_{\theta}(\tau|y_{2})$ that minimizes the expected transportation cost. Nonetheless, we can choose a certain coupling $\gamma\in\Gamma(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ to obtain an upper bound of this Wasserstein term, i.e.,
@@
Moreover, if we fix the predicted variance $\sigma_{\theta}$ of the planner as a fixed constant, then the expression of $\hat{W}_{1}(y_{1},y_{2};\theta)$ can be further simplified to $\hat{W}_{1}(y_{1},y_{2};\theta)=\sum_{t}\|\mu_{\theta}(s^{1}_{1:t},y_{1},t)-\mu_{\theta}(s^{2}_{1:t},y_{2},t)\|$. The overall AIGB-Pearl algorithm is summarized in Algorithm 1 in Appendix D due to page limits.
4 Experiments
We conduct both simulated and real-world experiments to validate the effectiveness of our approach. In the experiments, we mainly investigate the following Research Questions (RQs): (1) Does enhancing AIGB with policy optimization improve overall performance, and can it generalize better to unseen data compared to existing AIGB methods? (Section 4.2) (2) How does the KL-Lipschitz constraint affect the performance of the planner? (Section 4.3) (3) Can the proposed method guarantee the Lipschitz property of the evaluator and the planner? (Section 4.4). The training stability of AIGB-Pearl is studied in Appendix E.5.
@@ 4 Experiments
4.1 Experiment Setup
Baselines. We compare our method with the state-of-the-art AIGB methods, including DiffBid (Guo et al., 2024) and DT (Chen et al., 2021a) that learn from conditional behavior cloning of the offline dataset with a diffusion model and a Causal Transformer, respectively. We also compare our method with RL auto-bidding methods, including USCB (He et al., 2021) that learns the auto-bidding policy in a manually constructed advertising system with DDPG (Silver et al., 2014); and offline RL auto-bidding methods, including model-free offline RL methods BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and Diff-QL (Wang et al., 2022), and model-based offline RL method MOPO (Yu et al., 2020).
4.2 Overall Performance

### v3

@@ Abstract
Auto-bidding serves as a critical tool for advertisers to improve their advertising performance. Recent progress has demonstrated that AI-Generated Bidding (AIGB), which learns a conditional generative planner from offline data, achieves superior performance compared to typical offline reinforcement learning (RL)-based auto-bidding methods. However, existing AIGB methods still face a performance bottleneck due to their inherent inability to explore beyond the static offline dataset. To address this, we propose AIGB-Pearl (Planning with EvaluAtor via RL), a novel method that integrates generative planning and policy optimization. The core of AIGB-Pearl lies in constructing a trajectory evaluator for scoring generation quality and designing a provably sound KL-Lipschitz-constrained score maximization scheme to ensure safe and efficient exploration beyond the offline dataset. A practical algorithm incorporating the synchronous coupling technique is further devised to ensure the model regularity required by the proposed scheme. Extensive experiments on both simulated and real-world advertising systems demonstrate the state-of-the-art performance of our approach.
@@ 1 Introduction
Hence, there arises a question: built on AIGB, the latest state-of-the-art auto-bidding method, can we devise a plausible scheme to involve policy optimization in its generative model? To this end, a natural idea is to integrate offline RL methods into AIGB. However, it is nontrivial to implement in the auto-bidding problem since (i) there is a lack of reward signals in AIGB to guide the generative model. Specifically, the generation quality of the generative model remains unknown during training, making it infeasible to explore new trajectories beyond the offline dataset; (ii) no dedicated offline RL algorithm exists for AIGB. In particular, theoretical approaches to ensure safe generalization and mitigate OOD issues for generative models in auto-bidding remain largely unexplored.
To address these critical challenges, we propose AIGB-Pearl (Planning with EvaluAtor via RL), an RL-enhanced version of AIGB that learns a trajectory evaluator to score generation quality and drive exploration of the generative model through continuous interaction. The evaluator is trained through supervised learning on the offline dataset. Crucially, to mitigate the OOD problem and ensure reliable evaluator utilization, we examine the theoretical upper bound on the evaluator’s bias. Then, guided by this analysis, we establish a KL-Lipschitz-constrained score maximization objective with a provable sub-optimality bound, enabling safe and effective exploration beyond the offline data. Moreover, to perform constrained score maximization, we design a practical algorithm incorporating the synchronous coupling technique, which helps ensure the Lipschitz requirement of the generative model. In addition, we note that AIGB-Pearl operates without the need for bootstrapping, exhibiting greater training stability compared to offline RL methods.
To summarize, our contributions in this paper are fourfold: (i) we propose a novel generative auto-bidding method, AIGB-Pearl, that enables continuous improvement in generation quality through exploration beyond the offline dataset; (ii) we propose a provable KL-Lipschitz constrained score maximization objective with a sub-optimality bound, ensuring a safe and effective generalization beyond the offline dataset; (iii) we devise a practical algorithm with synchronous coupling that effectively ensures the Lipschitz requirement for the generative model; (iv) extensive simulated and real-world experiments demonstrate that AIGB-Pearl achieves SOTA performance and verify the effectiveness of the developed techniques in enhancing safe and effective generalization.
@@ 2.2 Offline RL Methods
Due to the offline setting of the considered auto-bidding problem, directly employing Eq. 2 results in the infamous out-of-distribution (OOD) problem (Fujimoto et al., 2019), making the policy erroneously deviate from the offline dataset $\mathcal{D}$. As a standard solution, offline RL (Yu et al., 2020; Kumar et al., 2020; Kidambi et al., 2020; Wang et al., 2022b) constrains the policy’s behavior near $\mathcal{D}$ during TD learning, enabling reliable generalization beyond the offline dataset.
@@ 2.3 Generative Auto-bidding Methods
Unlike RL methods, the AI-generated auto-bidding (AIGB) (Guo et al., 2024) treats the auto-bidding problem as a sequence generation task. Specifically, a conditional generative model is employed to fit the conditional trajectory distribution $p_{\theta}(\tau|y(\tau))$ within the offline dataset $\mathcal{D}$, i.e.,
@@
where $\theta$ denotes the parameter. Let $y_{m}>0$ be the maximum trajectory quality in $\mathcal{D}$, we have $\forall y\in\mathcal{D},y\in[0,y_{m}]$. During the inference stage, AIGB follows a planning-and-control architecture. Specifically, at each time step, a trajectory is sampled from the generative model, acting as the planner, with a manually set condition $y^{*}\triangleq(1+\epsilon)y_{m}$, where $\epsilon>0$ is a hyper-parameter with usually a small value. Then, an extra off-the-shelf inverse dynamic model (Agrawal et al., 2016), acting as the controller, is employed to derive the action. See Appendix B for detailed descriptions. AIGB avoids TD learning and generally outperforms offline RL methods (Guo et al., 2024).
@@ 3 Method
Enabling AIGB to explore higher-quality trajectories actively beyond the offline dataset can enhance its performance and generalization ability. To this end, we propose AIGB-Pearl (Planning with EvaluAtor via RL) that constructs a trajectory evaluator (referred to as the evaluator for simplicity) to integrate RL methods into AIGB’s planner training. Specifically, the evaluator learns a score $\hat{y}_{\phi}(\tau)$ to estimate the trajectory quality $y(\tau)$ via supervised learning, i.e., $\min_{\phi}\mathbb{E}_{\tau\sim\mathcal{D}}[(\hat{y}_{\phi}(\tau)-y(\tau))^{2}]$, where $\phi$ denotes the evaluator parameter. Then, with $\phi$ fixed, the planner tries to maximize the score of its generation through iterative interactions with the evaluator, as shown in Fig. 1. Formally, this can be formulated as:
@@
As can be observed, the effectiveness of AIGB-Pearl fundamentally hinges on the reliability of the evaluator. However, given the offline nature of the considered auto-bidding problem, the evaluator learning is restricted to the fixed dataset $\mathcal{D}$. As a result, directly solving Eq. 4 can induce the infamous OOD problem due to the evaluator’s generalization limits, potentially degrading the planner’s true performance. Crucially, there remains a lack of theoretical approaches to the OOD problem in RL-based training of the generative model.
To address this challenge, we examine the theoretical bounds on the evaluator’s bias. Then, guided by this analysis, we propose a KL-Lipschitz-constrained score maximization objective for the planner in Section 3.1 to ensure reliable utilization of the evaluator. Notably, this objective is theoretically justified by a sub-optimality bound established in Section 3.1.1. Finally, a practical algorithm is presented in Section 3.2, where tailored regularization methods are introduced to enforce Lipschitz continuity in both the evaluator and the planner.
@@ 3.1 KL-Lipschitz-constrained score maximization
Our basic idea is to optimize $\theta$ within a domain where the gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)\triangleq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]$ is bounded by a small certifiable upper bound. This ensures the score maximization occurs only in regions where the evaluator is relatively reliable.
@@
Based on this property, we incorporate a $\sqrt{T}R_{m}$-Lipschitz continuity constraint into the evaluator’s learning (as described in Section 3.2.1), and denote the Lipschitz constant of the resulting evaluator as $k\sqrt{T}R_{m}$, where $k\geq 0$ reflects the violation degree. A tighter satisfaction of the Lipschitz constraint by the evaluator results in $k$ closer to $1$.
@@
Note that $\delta_{D}$ could be regulated to a small value via supervised training of the evaluator, and $k$ depends on the Lipschitz property of the resulting evaluator  [^2] Note that $k$ cannot approach zero without compromising $\delta_{D}$, as excessively small $k$ prevents the evaluator from fitting the offline dataset $\mathcal{D}$.. Consequently, bounding the evaluator bias in the planner’s performance requires constraining the following two factors:
@@
The proof is given in Appendix C.3. Therefore, we constrain the planner’s Lipschitz constant to a positive value $L_{p}>0$ to ensure the boundedness of the first Wasserstein term, where $L_{p}$ is a hyperparameter whose lower bound is provided later in Eq. 10.
@@
Intuitively, the KL and Lipschitz constraints collectively ensure the planner’s generation under condition $y^{*}$ remains within a certified neighborhood of the offline dataset $\mathcal{D}$. At the same time, the evaluator maintains high accuracy within this $\mathcal{D}$-proximal region as it is trained on $\mathcal{D}$, and the Lipschitz regularization on the evaluator bounds its sensitivity to input perturbations, preventing drastic value fluctuations in OOD regions and promoting more reliable extrapolation. Therefore, Eq. 8 with a Lipschitz-regularized evaluator can realize reliable evaluator-guided score maximization.
@@
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]$ be the expected distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$. The true performance gap between the optimal parameter $\theta^{*}$ and the solution $\hat{\theta}$ to Eq. 8 is bounded by:
@@
Theoretical Result Analysis. In Theorem 3, the constants $\delta_{M},R_{m},T,\epsilon,y_{m},\rho$ characterize domain-specific properties of the auto-bidding task and the offline dataset $\mathcal{D}$. Nonetheless, a lower training error $\delta_{D}$ of the evaluator and a closer $k$ to $1$ correspond to a smaller sub-optimality gap. Note that $k$ cannot approach zero without compromising $\delta_{D}$, as excessively small $k$ prevents the evaluator from fitting the offline dataset $\mathcal{D}$.
@@ 3.2.1 Lipschitz Trajectory Evaluator
where $\beta_{1}>0$ is a hyper-parameter, $\cdot]_{+}\triangleq\max\{0,\cdot\}$. Moreover, to further enhance the accuracy of the evaluator, we devise two specific techniques, including the LLM Embedding enhancement and pair-wise learning, whose details are given in Appendix D.1.
@@ 3.2.2 Lipschitz Planner With Synchronous Coupling
Wasserstein Upper Bound as Surrogate. Accurate calculation of this Wasserstein distance term is computationally challenging, as it requires finding the optimal coupling between $p_{\theta}(\tau|y_{1})$ and $p_{\theta}(\tau|y_{2})$ that minimizes the expected transportation cost. Nonetheless, we can choose a certain coupling $\gamma\in\Gamma(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ to obtain an upper bound of this Wasserstein term, i.e.,
@@
Moreover, if we fix the predicted variance $\sigma_{\theta}$ of the planner as a fixed constant, then the expression of $\hat{W}_{1}(y_{1},y_{2};\theta)$ can be further simplified to $\hat{W}_{1}(y_{1},y_{2};\theta)=\sum_{t}\|\mu_{\theta}(s^{1}_{1:t},y_{1},t)-\mu_{\theta}(s^{2}_{1:t},y_{2},t)\|$. The overall AIGB-Pearl is summarized in Algorithm 1 in Appendix D.2 due to page limits.
Remark 2.
We note that AIGB-Pearl can fully leverage the generalization capability of the conditional generative model. Specifically, the model acquires an implicit understanding of how trajectory characteristics vary with $y$ through conditional behavior cloning on $\mathcal{D}$. This enables the model not only to learn a rich and well-calibrated distribution over trajectories for conditions $y\in\mathcal{D}$, but also to generalize smoothly beyond the observed range, providing a high-quality initialization for OOD condition $y^{*}$. (In practice, we often initialize the search at $y^{*}$ using a pre-trained conditional generator as described in Algorithm 1.) Subsequently, under the guidance of reward maximization and constrained by Lipschitz continuity with respect to $y$, the model safely and efficiently explores improved trajectories without collapsing into unrealistic or unstable solutions.
@@
4 Experiments
We conduct both simulated and real-world experiments to validate the effectiveness of our approach. In the experiments, we mainly investigate the following Research Questions (RQs): (1) Does enhancing AIGB with policy optimization improve overall performance, and can it generalize better to unseen data compared to existing AIGB methods? (Section 4.2) (2) How does the KL-Lipschitz constraint affect the performance of the planner? (Section 4.3) (3) Can the proposed method guarantee the Lipschitz property of the evaluator and the planner? (Section 4.4). Moreover, the training stability of the planner is analyzed in Appendix E.5, and the accuracy of the evaluator is given in Appendix E.6, respectively.
4.1 Experiment Setup
Baselines. We compare our method with the state-of-the-art AIGB methods, including DiffBid (Guo et al., 2024) and DT (Chen et al., 2021a) that learn from conditional behavior cloning of the offline dataset with a diffusion model and a Causal Transformer, respectively. We also compare our method with RL auto-bidding methods, including USCB (He et al., 2021) that learns the auto-bidding policy in a manually constructed advertising system with DDPG (Silver et al., 2014); and offline RL auto-bidding methods, including model-free offline RL methods BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and Diff-QL (Wang et al., 2022b), and model-based offline RL method MOPO (Yu et al., 2020).